OpenStudy (mislovely187):

The sides of an isosceles triangle are 5, 5, and 7. Find the measure of the vertex angle to the nearest degree.

OpenStudy (denonakavro):

do you know law of cosines

OpenStudy (mislovely187):

c2=a2+b2-2(ab) cos C

OpenStudy (denonakavro):

49=50-50cosC 1=50cosC 1/50=cosC Then use a graphing calculator and plug this in cos^-1(1/50) = 88.854 degrees

OpenStudy (denonakavro):

If you do not have a calculator with you for later then you can use this awesome tool http://www.calculatorsoup.com/calculators/geometry-plane/triangle-law-of-cosines.php

OpenStudy (mislovely187):

thanks

OpenStudy (david27):

Hey

OpenStudy (kanwal32):

sin inverse 0.7